# DBCopilot: Scaling Natural Language Querying to Massive Databases

### v2

@@ 4.1.2. Baselines
We compare DBCopilot to various sparse/dense retrieval techniques for scaling NL querying over massive databases, including zero-shot, LLM-enhanced, and fine-tuned methods:
  • BM25 (Wikipedia contributors, 2023a): BM25 is a well-known ranking function widely used in search systems to retrieve relevant documents.
@@
  • CRUSH (Kothyari et al., 2023): CRUSH adopts a two-stage approach to retrieve schema elements for NL2SQL. It first instructs the LLM to generate a hallucinated schema, and then combines multiple retrieval results to obtain the actual schema.
  • DTR (Herzig et al., 2021): DTR is an approach that uses contrastive learning on labeled (question, table) pairs to develop table retrievers for open-domain question answering over tables.
@@ 4.1.3. Evaluation Metrics
For schema routing, we use Recall@k and mAP (mean average precision) to evaluate the relevance of retrieved databases and tables. Recall@k measures the fraction of relevant instances in the top-k retrieved candidates. mAP calculates the average precision over all queries, which reflects the overall retrieval performance. For SQL generation, traditional surface-level metrics are inappropriate for LLM-based approaches as they may produce SQL queries that differ greatly from the ground truth. Following recent studies (Gu et al., 2023b; Gao et al., 2023; Li et al., 2023b), we use EX (execution accuracy), which compares the results of generated SQL queries with corresponding ground truths. We also report the cost of LLM invocations.
@@ 4.1.4. Implementation Details
We implemented DBCopilot using Python 3.10, Pytorch 2.1, and the Transformers (Wolf et al., 2020) library.
Schema Routing.  We adopted T5-base (Raffel et al., 2020) as the backbone of the schema questioner and router. We set the batch size to $32$, the learning rate to $5\times 10^{-5}$, and the maximum number of training epochs to $20$ for all models. Model parameters were optimized using AdamW (Loshchilov and Hutter, 2019) and a linear learning rate schedule with no warm-up steps. We used the schema questioner trained on Spider and Bird to generate pseudo-questions for Fiben, which has no training data. For each database collection, we synthesized $10^{5}$ training instances.
SQL Generation.  We used gpt-3.5-turbo-16k-0613 to select the optimal schema and generated the SQL query with the OpenAI demo prompt (Gao et al., 2023). We set the temperature to 0 for reproducibility.
Baselines.  To maintain fairness, all baselines compared in this study relied on schema information only. We treated tables as retrieval targets, with the content of each table being the flat normalized names of the table and its columns. For each NL question, we retrieved the top 100 tables and determined database scores based on the average of their table scores. We used the standard Okapi BM25 with two adjustable parameters for BM25 and adopted the best performing all-mpnet-base-v2 for SXFMR. BM25 and DTR were fine-tuned on synthetic data, consistent with DBCopilot.
4.2. Schema Routing
We first compare the performance of schema routing, with the results shown in Table 3, Table 4, Figure 5, and the following findings.
Table 3 shows the database and table recall on regular test sets. We can see that DBCopilot consistently outperforms all types of baselines by a large margin. The database recall@1 of DBCopilot surpasses the second-best approach by up to 12.67% and 11.93% on the Spider and Bird datasets, respectively. This result verifies that DBCopilot can effectively identify the target database of an NL question. Furthermore, DBCopilot also holds an advantage in table recall, with a lead of up to 9.59%, which is crucial for providing complete tables for SQL generation. Even on the single database dataset Fiben, DBCopilot still achieves a 3.4% improvement. These results confirm that DBCopilot can effectively navigate NL questions to their target databases and tables over massive databases.
@@ 4.2. Schema Routing
Table 4. Schema routing performance on robustness test sets.
Method Spidersyn Spiderreal
@@
DBCopilot 62.67 85.11 70.35 86.26 79.53 95.47 87.93 96.06
When querying over massive databases, data consumers without schema knowledge typically express NL questions that are highly inconsistent with database schemas. To evaluate the robustness in this situation, Table 4 shows the results of Spidersyn and Spiderreal, which replace explicitly mentioned schema elements in NL questions posed by NL2SQL annotation. We can see that the performance of all methods decreases with this diversity setting. Sparse retrieval baselines are significantly affected as they rely heavily on the vocabulary matching assumption, and the LLM-enhanced method also struggles with this challenge, despite being designed specifically for this case. Compared to these baselines, DBCopilot is the least affected and therefore maintains a larger advantage with 14.89% and 19.88% on database recall@1, as well as 2.83% and 6.94% on table recall@5. We believe this is because by injecting schema knowledge through schema-specific training, our method can better understand the semantics of database schemas and effectively learn the mapping between questions and schemas.
@@
Finding 2. DBCopilot is more robust when natural language questions are highly inconsistent with database schemas.
@@
Figure 5. (a) Table mAP w.r.t. different numbers of database tables. (b) Table recall@$k$ w.r.t. the number of retrieved tables.
DBCopilot model schema routing as a relation-aware, generative joint retrieval process, which differs from traditional baselines of independent retrieval. To assess the effect of relation modeling and joint retrieval, we evaluate the performance under different numbers of database tables and retrieved tables. As shown in Figure 5 (a), the mAP of DBCopilot is relatively stable as the number of database tables increases, while the independent retrieval baselines decrease significantly. The recall of DBCopilot is also consistently higher than the baselines as the number of retrieved tables increases from Figure 5 (b). These results verify the scalability of DBCopilot through relation modeling and generative routing.
@@
Finding 3. Relation modeling and joint retrieval are essential for DBCopilot to perform effective schema routing.
4.3. Schema-Agnostic NL2SQL
@@
We then compare the schema-agnostic NL2SQL performance under gold schemas, different routing methods, and prompt strategies.
@@
Oracle Test.  We conduct oracle tests to access the upper bound of schema-agnostic NL2SQL, by providing gold schemas at various small scales. Specifically, we gradually move from 5 database schemas to the gold database, then to gold tables only, or even to gold tables with only gold columns. Experimental results in Table 5 show that as the number of schema elements increases, the EX of LLM-generated SQL queries continuously decreases, while the inference cost keeps increasing. This suggests that LLMs, like PLMs (Li et al., 2023c), are also affected by extraneous schema elements in SQL generation. Predictably, the situation gets worse as the scale of the schema increases. As a result, it is highly inefficient and ineffective to expose massive database schemas directly to LLMs.
@@
Table 5. Schema-agnostic NL2SQL performance with gold (G.) schemas, different routing methods, and prompt strategies.
Routing Method Provided Schema Prompt Strategy Spider Bird Spidersyn
@@
G. Column Basic 79.69 0.40 35.98 0.75 76.60 0.40
G. Table Basic 77.08 0.52 29.86 1.44 67.21 0.52
G. Database Basic 74.47 0.81 25.75 2.70 63.73 0.81
5 Database Basic 67.89 2.98 19.10 8.80 49.90 2.99
CRUSHBM25 Top 1 Basic 56.77 1.30 13.69 2.67 20.79 1.23
DTR Top 1 Basic 45.16 0.66 17.60 2.12 26.98 0.62
DBCopilot Top 1 Basic 64.12 0.69 21.97 2.35 37.81 0.70
DBCopilot Top 5 Basic 61.51 2.28 19.56 7.18 34.53 2.27
DBCopilot Top 5 COT 57.83 5.33 19.17 16.49 33.95 5.29
@@
Finding 4. Schema routing is necessary for LLM-based schema-agnostic NL2SQL for both efficiency and effectiveness.
Impact of Improved Recall.  We perform schema-agnostic NL2SQL to verify the impact of improved recall on the accuracy of generated SQL queries. Without loss of generality, we compare DBCopilot to CRUSHBM25 (sparse) and DTR (dense), which achieve suboptimal routing performance on Spider and Bird, respectively. As shown in Table 5, DBCopilot outperforms these two baselines by up to 10.83% EX. The LLM inference cost of DBCopilot is slightly higher, possibly due to successfully retrieving more required tables. We can also confirm the correlation between routing recall and EX by comparing these two baselines. CRUSHBM25 outperforms DTR in recall and EX on Spider, while DTR achieves higher EX on the other two datasets due to its superior recall. Therefore, the success of schema routing is essential for schema-agnostic NL2SQL.
Finding 5. DBCopilot outperforms retrieval baselines in schema-agnostic NL2SQL due to its superior schema routing.
Exploration of Prompt Strategies.  By generating multiple candidate schemas, DBCopilot can effectively increase the recall of the target schema. We explore three prompt strategies for leveraging multiple schemas. Interestingly, SQL generation with the top 1 schema performs best, even though the top 5 candidates have 22.44% more database recall on Spidersyn than the top 1. This may be due to confusion caused by similar candidate schemas, thus requiring more database/human information to disambiguate. By introducing human-in-the-loop (HITL) to select the desired schema from the top 5 candidates, DBCopilot comes close to the ideal situation of providing the gold database schema, but with lower inference costs. This highlights the ability of DBCopilot to exclude extraneous tables and its promise for schema-agnostic NL2SQL.
Finding 6. Using the best schema prompting strategy for SQL generation is optimal unless there is human interaction.
4.4. Ablation Study
Table 6. Performance loss in ablation studies. “BS”, “OD”, “CD”, “DB” stands for basic serialization, original NL2SQL training data, constrained decoding, and diverse beam search.
Spider Bird
@@ 4.4. Ablation Study
   w/ BS -3.20 -2.90 -4.09 -5.75 -3.39 -1.63 -6.43 -5.99
   w/ OD -84.72 -82.40 -88.76 -79.55 -88.92 -82.27 -85.73 -88.81
   w/o CD -0.39 -0.19 -4.15 -2.51 +0.12 -0.45 -2.81 -1.93
   w/o DB +0.48 -3.00 +0.64 -2.88 -0.52 -4.89 -3.69 -8.87
We finally perform ablation studies on DBCopilot to validate the contribution of each module, as summarized in Table 6.
DFS Serialization.  DFS serialization is essential for schema routers to learn the underlying relations of databases and tables. To demonstrate this, we replace the DFS serialization with the unordered basic serialization of tables, similar to DSI for multiple documents. As shown in Table 6, this results in a 1% to 6% reduction in recall. DFS serialization effectively reduces the search space, whether by first generating a database and then generating its tables, or by continuing to generate based on existing tables. This graph-based serialization also makes constrained decoding possible.
Data Synthesis.  Generative retrieval is difficult to generalize to unseen documents during training. As we can see, routers trained on the original training instances ($N$, $S$) extracted from NL2SQL datasets are almost completely incapable of routing on test schemas, even when constrained decoding is used. Data synthesis by sampling schemas and generating pseudo-questions can effectively solve this challenge. Figure 6 shows how the performance varies with the amount of synthetic data. Database and table recalls increase rapidly at first as more data is synthesized, but level off after a certain amount. How to improve the quality of synthetic data and reduce the need for it may be a future direction.
Decoding Strategy.  Decoding optimizations (i.e., constrained decoding and group beam search) are beneficial for generative schema routing during inference. On the one hand, constrained decoding is useful for generating accurate table identifiers, since the drop from removing it is mainly in table recalls. On the other hand, group beam search helps to generate multiple diverse schemas because database recall@5 decreases significantly without it.
Finding 7. DFS serialization, data synthesis, and decoding optimizations contribute to DBCopilot in different ways.
5. Related Work
5.1. Natural Language to SQL
NL2SQL is an essential part of the natural language interface. Most research has approached this task along the lines of semantic parsing (Kamath and Das, 2019; Li et al., 2020). Early NL2SQL parsing systems adopted symbolic approaches that relied heavily on rule or grammar engineering, making them difficult to apply widely (Zelle and Mooney, 1996; Woods, 1973). In recent years, with the advances in deep learning, especially PLMs, end-to-end NL2SQL approaches that using neural generation models have become the mainstream (Li et al., 2023c; Wang et al., 2020; Gu et al., 2023a; Li et al., 2023a; Yu et al., 2021). Some efforts have been made to improve encoding and decoding, such as relation-aware encoding (Wang et al., 2020; Li et al., 2023a) and grammar-based decoding (Xu et al., 2017; Scholak et al., 2021). Some other efforts have focused on generating accurate SQL queries through schema linking (Li et al., 2023c; Wang et al., 2020) or skeleton generation (Gu et al., 2023a, b).
LLMs have emerged as a new paradigm for NL2SQL (Rajkumar et al., 2022; Liu et al., 2023a; Gao et al., 2023). Unlike previous work, the core of the LLM-based NL2SQL solution is prompt engineering, i.e., prompting LLM to generate correct and efficient SQL queries. Several efforts have been made to select appropriate in-context learning examples through question classification (Pourreza and Rafiei, 2023), skeleton retrieval (Guo et al., 2023), and in-domain example synthesis (Chang and Fosler-Lussier, 2023). Research has also been done on decoupling zero-shot NL2SQL into PLM-based sketch generation and LLM-based SQL completion (Gu et al., 2023b). These studies are orthogonal to our work and can be easily integrated to optimize DBCopilot SQL generation.
5.2. Table Discovery
Table discovery techniques are critical to data management, allowing users to explore and gain insights from data warehouses or lakes (Fan et al., 2023a). A common practice for table discovery is query-driven discovery (Zhang and Ives, 2020; Nargesian et al., 2019), where the query is typically in the form of keywords, tables, or NL questions. For keyword search, data discovery systems find relevant tables by searching for topic terms in the metadata (Cafarella et al., 2009; Pimplikar and Sarawagi, 2012). For table search, tables are retrieved to augment a given table as the query (Sarma et al., 2012). Joinable table search (Dong et al., 2021; Zhu et al., 2019) and unionable table search (Fan et al., 2023b; Khatiwada et al., 2023) aim to augment a query table with additional attributes and new tuples, respectively. For NL question search, open domain question answering retrieves independent tables to answer the question with table-based models (Chen et al., 2021; Herzig et al., 2021). Although dense retrieval methods based on contrastive learning have achieved advanced performance in the above tasks (Dong et al., 2021; Fan et al., 2023b; Herzig et al., 2021), they are difficult to consider the structural relationships between tables and retrieve them jointly. Alternatively, data navigation allows users to navigate interactively through the tables organized hierarchically (Nargesian et al., 2020) or relationally (Fernandez et al., 2018). Our work can be seen as automated data navigation, where we employ DSI to model the query-based navigation in an end-to-end manner.
6. Conclusion
In this paper, we present DBCopilot, a framework for effectively scaling natural language querying to massive databases. DBCopilot decouples schema-agnostic NL2SQL into schema routing and SQL generation through LLM-Copilot collaboration. We present an approach to effectively navigate the target database and tables for an NL question in a relation-aware, end-to-end manner. We also propose a training data synthesis paradigm to adapt routers over massive databases without manual intervention, and explore various prompt strategies to incorporate multiple candidate schemas. Extensive experiments verify the effectiveness of our solution. In summary, DBCopilot pushes the boundaries of NL2SQL, enabling researchers to better strategize their quest for data accessibility.
Appendix A Supplementary Experiments
A.1. Case Study

### v3

@@ 4.1.3. Baselines
We compare DBCopilot to various sparse and dense retrieval techniques for schema routing (RQ1 & RQ2), including zero-shot, LLM-enhanced, and fine-tuned methods:
  • BM25 (Wikipedia contributors, 2023b): BM25 is a renowned ranking function that serves as a standard in search engines to rank documents by their relevance to a specified query.
@@
  • CRUSH (Kothyari et al., 2023): CRUSH adopts a two-stage approach to retrieve schema elements for NL2SQL. It first instructs an LLM to generate a hallucinated schema, and then combines multiple retrieval results and reranks them based on their relationships to obtain the actual schema.
  • DTR (Herzig et al., 2021): DTR is an approach that uses contrastive learning on (question, table) pairs to develop table retrievers for open-domain question answering over tables.
For schema-agnostic NL2SQL (RQ3 & RQ4), since existing NL2SQL methods cannot be directly applied to this problem even when equipped with schema linking (as discussed in § 2.1), we combine a representative LLM-based NL2SQL method (Gao et al., 2024) with different retrieval methods for comparison. This allows us to isolate and evaluate the effects of schema routing and prompt strategies on end-to-end SQL generation in a controlled setting.
@@ 4.1.4. Evaluation Metrics
Different evaluation metrics are used for schema routing and SQL generation respectively.
For schema routing, we use Recall@k and mAP (mean average precision) to evaluate the relevance of retrieved databases and tables. Recall@k measures the fraction of relevant instances in the top-k retrieved candidates. mAP calculates the average precision over all queries, which reflects the overall retrieval performance.
For SQL generation, traditional surface-level metrics are inappropriate for LLM-based approaches as they may produce SQL queries that differ greatly from the ground truth. Following recent studies (Gu et al., 2023b; Gao et al., 2024; Li et al., 2023b), we use EX (execution accuracy), which compares the results of generated SQL queries with corresponding ground truths. We also report the cost (\$) of LLM invocations.
@@ 4.1.5. Implementation Details
We implemented DBCopilot using Python 3.10, Pytorch 2.1, along with some major libraries such as Transformers (Wolf et al., 2020) and Pytorch Lightning (Falcon and The PyTorch Lightning team, 2019). The source code is available at github.com/tshu-w/DBCopilot.
Schema Routing.  We adopted T5-base (Raffel et al., 2020) as the backbone of the schema questioner and router. We trained a unified schema questioner on Spider and Bird NL2SQL training sets and three schema routers on each database collection (Spider, Bird, and Fiben). For each database collection, we sampled $1\times 10^{5}$ SQL query schemata uniformly from the constructed graph, covering all (100%) databases and tables, and synthesized the corresponding training instances using the schema questioner. We set the batch size to $32$, the learning rate to $5\times 10^{-5}$, and the maximum number of training epochs to $20$ for all models. Model parameters were optimized using AdamW (Loshchilov and Hutter, 2019) and a linear learning rate schedule with no warm-up steps. To detect joinable tables besides primary and foreign, we adopted a heuristic way that two tables are joinable if the exact match overlap (Jaccard similarity) of their column values is greater than 0.85. During schema routing, we generated 10 schema sequences for each question using diverse beam search, with beams and beam groups set to 10 and the diversity penalty set to 2.0.
SQL Generation.  We used gpt-3.5-turbo-0125222https://platform.openai.com/docs/models#gpt-3-5-turbo to select the optimal schema and generated the SQL query, with the generation temperature set to 0 for reproducibility.
Baselines.  To maintain fairness, all baselines compared in this study relied on schema information only. We treated tables as retrieval targets, with the content of each table being the flat normalized names of the table and its columns. For each NL question, we retrieved the top 100 tables and ranked the databases based on the average score of their retrieved tables. We used the standard Okapi BM25 with two adjustable parameters for BM25 and adopted the best performing all-mpnet-base-v2 for SXFMR. BM25 and DTR were fine-tuned on synthetic data consistent with DBCopilot to verify our relation-aware joint retrieval.
All experiments were run using Docker image nvidia/cuda:12.1.1-cudnn8-devel-ubuntu22.04 on a server with AMD EPYC 7H12 64-Core Processors, 1TB RAM, and NVIDIA A100 40G GPUs.
4.2. Schema Routing (RQ1 & RQ2)
4.2.1. Experimental Results
We first compare the performance of schema routing on both regular and robustness datasets, with the experimental results shown in § 4.1.2, § 4.2.1, Figure 7, and the following findings.
§ 4.1.2 shows the database and table recall on regular test sets. We can see that DBCopilot consistently outperforms all zero-shot, LLM-enhanced, or fine-tuned baselines by a large margin. The database recall@1 of DBCopilot surpasses the second-best approach by up to 12.67% and 11.93% on the Spider and Bird datasets, respectively. This result verifies that DBCopilot can effectively identify the target database of an NL question. Furthermore, DBCopilot also holds an advantage in table recall, with a lead of up to 9.59%, which is crucial for providing complete tables for SQL generation. Even on the single database dataset Fiben, DBCopilot still achieves a 3.4% improvement. These results confirm that DBCopilot can effectively navigate NL questions to their target databases and tables over massive databases.
@@ 4.2.1. Experimental Results
When querying over massive databases, data consumers without schema knowledge typically express NL questions that are highly inconsistent with database schemata. To evaluate the robustness of different methods in this situation, § 4.2.1 shows the experimental results on Spidersyn and Spiderreal, which replace explicitly mentioned schema elements in NL questions posed by NL2SQL annotation. We can see that the performance of all methods decreases in this actual scenario setting. Sparse retrieval baselines are significantly affected as they rely heavily on the vocabulary matching, and the LLM-enhanced method also struggles with this challenge, despite being designed specifically for this case. Compared to these baselines, DBCopilot is the least affected and therefore maintains a larger advantage with 14.89% and 19.88% on database recall@1, as well as 2.83% and 6.94% on table recall@5. We believe that by injecting schema knowledge through schema-specific training, DBCopilot can better understand the semantics of database schemata and effectively learn the semantic mapping between NL questions and schemata.
Table 4. Schema routing performance on robustness tests.
{NiceTabular}
[tabularnote = We use the same indicators and markups as § 4.1.2.]@ccccccccc@ Method Spidersyn Spiderreal
@@
DBCopilot 62.67♯‡ 85.11 70.35♯‡ 86.26♯‡ 79.53♯‡ 95.47♯‡ 87.93♯‡ 96.06♯‡
@@
Finding 2. DBCopilot is more robust when NL questions are highly inconsistent with database schemata.
@@
Figure 7. (a) Table mAP w.r.t. different numbers of database tables. (b) Table recall@k w.r.t. different numbers of retrieved tables.
We further analyzed the performance regarding different database sizes (i.e., the number of tables contained) and the number of tables retrieved. Experimental results are shown in Figure 7 (a) and (b), respectively. First, we categorize the instances based on the number of tables contained in their target database and compare the table mAP across different database sizes. As shown in the Figure 7 (a), the mAP of DBCopilot is relatively stable as the number of database tables increases, while the independent retrieval baselines decrease significantly. We believe this is because by modeling schema routing as a relation-aware joint retrieval process, DBCopilot can better capture the relationships within the database compared to independent retrieval methods, which is critical for routing databases that contain many tables. Furthermore, as shown in Figure 7 (b), the table recall of DBCopilot is consistently higher than that of the baselines as the number of retrieved tables increases, achieving a recall@20 that exceeds the recall@50 of all baselines. This suggests that retrieving more tables and reranking them based on their relationships can hardly compensate for the advantages of our end-to-end relation-aware joint retrieval approach.
@@
Finding 3. Relation-aware joint retrieval are essential for DBCopilot to perform effective schema routing.
4.2.2. Case Study
[…]
Figure 8. Successful case of DBCopilot in schema routing.
We provide successful cases to illustrate how DBCopilot improves schema routing through schema-specific training and joint retrieval. As shown in Figure 8, DBCopilot is the only approach that accurately routes the appropriate schema for the given question. Sparse retrieval methods (i.e., BM25 and CRUSHBM25) rely heavily on lexical matching while ignoring the semantic matching between the question and the schemata. However, untuned dense retrieval methods (i.e., SXFMR and CRUSHSXFMR) also fail to capture the underlying semantics of tables and retrieve the wrong databases. Fine-tuning based approaches identify the correct database where DTR misses the country table linked to the countrylanguage table, but our relation-aware DBCopilot does not.
While DBCopilot consistently outperforms the baselines, there are a few cases where DTR retrieves the correct schema but DBCopilot does not. For example, Figure 9 shows that DBCopilot confuses the airlines and airports tables of database flight, but DTR covers the correct tables by retrieving more loosely. Upon further analysis, we find that the quality of synthetic data impacts the robustness of schema routing. This is attributed to two main factors: 1) hallucination (Papicchio et al., 2023), where the generated NL questions deviate from sampled schemata; 2) generation bias, where the simple pipeline lacks diversity and quality control mechanisms. Developing more sophisticated data synthesis techniques to address these challenges remains an important direction for future work.
[…]
Table 5. Method efficiency & resource consumption.
[…]
BM25 1677.1 10 36 800 -
SXFMR 49.0 48 420 2,230 3742
[…]
BM25 1677.1 6972 36 800 -
DTR 49.0 5161 420 2,230 3742
DBCopilot 43.0 13678 852 656 11807
4.2.3. Efficiency & Resource Consumption
We also measure the efficiency and resource consumption of each schema routing method. § 4.2.2 shows the number of queries per second (QPS), the duration of build time, the amount of disk space used, and the consumption of system and GPU memory. LLM inference and complex post-processing are the main reasons for the slowness of CRUSH approaches. DBCopilot, based on an encoder-decoder architecture, consumes more disk space and GPU memory than dense retrieval approaches that based on an encoder-only architecture. Despite the slightly higher resource consumption for training and inference in generative retrieval, we believe and have observed that the growing popularity of generative AI is leading to an increasing number of proposed optimizations.
4.3. Schema-Agnostic NL2SQL (RQ3 & RQ4)
@@
This section compares the performance of schema-agnostic NL2SQL under various situations, including providing gold schemata, using different schema routing methods, and prompt strategies. Experimental results are shown in § 4.3.
@@
Oracle Test.  We first conduct oracle tests to access the upper bound of schema-agnostic NL2SQL, as well as the impact of irrelevant schema elements for LLM-based NL2SQL, by providing gold schemata at various small scales. Specifically, we gradually move from 5 database schemata to the gold database, then to gold tables only, or even to gold tables with only gold columns. Experimental results in § 4.3 show that as the number of schema elements increases, the EX of LLM-generated SQL queries continues to decrease significantly, while the inference cost keeps increasing many times over. This suggests that LLMs, like PLMs (Li et al., 2023c), are also affected by extraneous schema elements in SQL generation. Predictably, the situation gets worse as the schema scales. As a result, it is highly inefficient and ineffective to expose massive database schemata directly to LLMs, making schema routing necessary for schema-agnostic NL2SQL.
@@
Table 6. Evaluation of schema-agnostic NL2SQL with various gold schemata, different schema routing methods, and prompt strategies.
[…]
notes/style=0 {NiceTabular}[tabularnote = “T.” and “C.” stand for table and column, respectively. We bold best results of schema-agnostic NL2SQL and underline the best with human in the loop to select the optimal candidate schema. For retrieval baselines, a candidate schema consists of the top candidate database and the retrieved tables that belong to the database. ]@cccccccc@ Routing Method Prompt Template Spider Bird Spidersyn
@@
Oracle Test
Gold T. & C. 81.24 0.07 35.98 0.15 77.37 0.07
Gold T. 76.60 0.09 31.03 0.24 68.67 0.09
Gold DB 74.08 0.12 27.57 0.40 61.99 0.13
5 DB w. Gold 70.21 0.13 18.38 8.80 51.64 0.40
Best Schema Prompting
CRUSHBM25 Top 1 56.58 0.10 14.73 0.26 20.60 0.09
DTR Top 1 43.91 0.11 19.04 0.33 26.21 0.10
DBCopilot Top 1 63.35 0.11 23.47 0.36 37.43 0.11
Multiple Schema Prompting
CRUSHBM25 Top 5 55.61 0.20 17.60 0.57 21.86 0.19
DTR Top 5 54.45 0.27 19.56 0.79 32.59 0.27
DBCopilot Top 5 61.51 0.31 21.45 0.95 35.49 0.31
Multiple Schema COT Prompting
CRUSHBM25 Top 5 53.97 0.28 22.10 0.83 22.34 0.26
DTR Top 5 50.68 0.36 22.43 1.08 28.53 0.37
DBCopilot Top 5 51.26 0.40 22.75 1.24 31.24 0.39
Human in the Loop
CRUSHBM25 Top 5 69.34 0.10 25.10 0.32 34.72 0.09
DTR Top 5 66.92 0.11 25.42 0.37 47.78 0.11
@@
Finding 4. Schema routing is necessary for scaling natural language querying to massive databases.
Impact of Improved Recall.  We then perform schema-agnostic NL2SQL to verify the impact of improved recall on the accuracy of generated SQL queries. Without loss of generality, we compare DBCopilot to sparse CRUSHBM25 and dense DTR baselines, which achieve suboptimal routing performance on Spider and Bird, respectively. As shown in § 4.3, DBCopilot outperforms CRUSHBM25 and DTR by 4.43%~11.22% EX when using the best schema prompting strategy and by 0.85%~20.07% EX when employing human in the loop to select the optimal schema from candidates. We compare and discuss the prompting strategies in the next paragraph. The LLM inference cost of DBCopilot is slightly higher, possibly due to its successful retrieval of more required tables. The correlation between schema routing recall and SQL generation EX can also be confirmed by comparing these two baselines. CRUSHBM25 outperforms DTR in recall and EX on Spider, while DTR achieves higher EX on the other two datasets due to its superior recall. Thus, the success of schema routing is essential for schema-agnostic NL2SQL.
Finding 5. DBCopilot leads in schema-agnostic NL2SQL due to its superior schema routing performance.
Exploration of Prompt Strategies.  By routing multiple candidate schemata, the recall of the target schemata for NL questions can be greatly improved. We explore three prompt strategies for leveraging multiple candidate schemata during SQL generation. As shown in § 4.3, when schema routing through DBCopilot, the performance of multiple schema prompting is slightly lower than that of single schema prompting. However, when schema routing through baselines, the performance of multiple schema prompting is typically better than that of single schema prompting. This is due to the trade-off between the improved recall brought by multiple candidate schemata and the impact of extraneous schemata on LLMs. As for multiple schema COT prompting, this strategy shows no absolute advantage over multiple schema prompting, and each wins and loses on different routing methods and datasets. Thus, the limitations of LLMs in distinguishing domain-specific schemata make it cost effective for DBCopilot to find the correct schema in the top 1 candidate. Finally, by introducing human-in-the-loop to select the desired schema from top 5 candidates, DBCopilot comes close to the ideal performance of providing the gold database schema, but with lower costs.
Finding 6. The best schema prompting strategy is optimal for DBCopilot unless there is human interaction.
4.4. Ablation Study (RQ5)
Table 7. Performance loss in ablation studies.
[…]
notes/style=0 {NiceTabular}@lcccccccc@ Spider Bird